OpenStudy (anonymous):

find the solution of the equation for x = 6. 8. y = 3x – 4

OpenStudy (anonymous):

y = 3(3) - 4 y = 9 - 4 y = 5

OpenStudy (anonymous):

find the solution of the equation for x = 6. 8. y = 3x – 4 (1 point) (6, 1) (6, 22) (6, 18) (6, 14)